OpenStudy (anonymous):

theta, given cos 2 theta =3/5 and theta terminates in quadrant I

terenzreignz (terenzreignz):

Half angle identities is all it takes ... \[\Large \cos\left(\frac \alpha 2\right)= \pm\sqrt{\frac{1+\cos(\alpha)}{2}}\]

OpenStudy (anonymous):

use identities ti find the exact values of sine and cosine

terenzreignz (terenzreignz):

Yes, well, here's a little something to get you started : \[\Large \cos(\theta) = \cos\left(\frac{2\theta}2\right)= \pm\sqrt{\frac{1+\color{red}{\cos(2\theta)}}2}\]

terenzreignz (terenzreignz):

And... you are, in fact, GIVEN the value of \(\large \color{red}{\cos(2\theta)}\)

OpenStudy (anonymous):